OpenStudy (anonymous):

Solve. x^2 + 8x = 24 answers: a. -4 +-2 sqrt 2 b. -4 +-2 sqrt 10 c. -4 +- 2 sqrt 5 d. -4 +- 2 sqrt 7 help pleasee :)

OpenStudy (anonymous):

x^2+8x-24=0 x^2+8x-24+16-16=0 (x+4)^2-24-16=0 (x+4)^2-40=0 follow that.

OpenStudy (anonymous):

soo what would the answer be ? @ amoodarya

OpenStudy (amoodarya):

x=-4 +-sqrt40 if you simplify that b is correct

OpenStudy (anonymous):

thanks so much